Exercise 5A-3 Cost Behavior; High-Low Method [LO5-10]

Hoi Chong Transport, Ltd., operates a fleet of delivery trucks in Singapore. The company has determined that if a truck is driven 153,000 kilometers during a year, the average operating cost is 12.5 cents per kilometer. If a truck is driven only 102,000 kilometers during a year, the average operating cost increases to 15.4 cents per kilometer.

Required:

1. Using the high-low method, estimate the variable operating cost per kilometer and the annual fixed operating cost associated with the fleet of trucks.

2. Express the variable and fixed costs in the form Y = a + bX.

3. If a truck were driven 127,000 kilometers during a year, what total operating cost would you expect to be incurred?
Answer #1
1
Kilometers Total cost
High level of activity 153000 19125 =153000*0.125
Low level of activity 102000 15708 =102000*0.154
Change 51000 3417
Variable cost per kilometer 0.067 =3417/51000
Fixed cost per year 8874 =19125-(153000*0.067)
2
Y = 8874 + 0.067 X
3
Total operating cost 17383 =8874+(127000*0.067)